The half-life of a substance is 120 days. The initial quantity is 72g.
a. Write the specific equation in the form A=Aoekt

b. How much will be left after 85 days?

c. How long until only 8 g is left?
1
Expert's answer
2011-07-20T11:42:48-0400
A=A0e-kt.
1/2 A0= A0e-kτ
Where τ is a half-life period. So we get
K=ln(2)/ τ
We have
A=A0e- ln(2)t/ τ = 72e- ln(2)t/120.
b.
A= 72e- ln(2)85/120 = 44.17 g.
c.
8= 72e- ln(2)t/120 -> 1/9=72e- ln(2)t/120 -> -2ln(3)= - ln(2)t/120
240 ln(3)/ln(2) = 379 days.